# Three-dimensional Abelian and nonAbelian gauge Higgs theories

### v1

@@ 5.3.3 The Coulomb-Higgs transition line
At continuous CH transitions scalar, charged, and gauge modes are all critical. As an example, in Fig. 6 we report numerical results obtained by FSS analyses (see A.3 for some details) taken from Refs. [87, 91], that provide evidence of a continuous transition for $N=25$ with critical scalar and charged modes. For instance, the left panel of Fig. 6 shows some data for the correlation length $\xi_{Q}$ computed in terms of the correlation function defined in Eq. (58), while the right panel shows the Binder parameter of the charged operator versus $\xi_{Q}/L$, computed in the Lorenz gauge in which it is equivalent to the Binder parameter of the scalar field, see Sec. 5.1.2. The excellent quality of the collapse of the data confirms that the transition is continuous and that correlations of both $Q_{\bm{x}}$ and ${\bm{\Gamma}}_{\bm{x}}$ are critical. Similar results have been obtained for the gauge correlation function.
@@
In Fig. 7 the numerical estimates of the exponent $\nu$ and $\eta_{q}$ are compared with the $1/N$ expressions reported in Sec. 4.1.3. The agreement is very good, thus providing a stringent check that the AHFT provides the correct effective description of the continuous transitions along the CH line. The same check has been performed for the exponent $\eta_{\Gamma}$, which has been determined by analyzing the susceptibility $\chi_{z}$ of the scalar field ${\bm{z}}_{\bm{x}}$ computed in the Lorenz gauge. Ref. [91] estimated $\eta_{\Gamma}$ for $N=25$, which, in turn provides an estimate of the RG dimension $y_{\Gamma}=(1+\eta_{\Gamma})/2$ of ${\bm{\Gamma}}_{\bm{x}}$, The result $y_{\Gamma}=0.4655(5)$ compares satisfactorily with the large-$N$ field-theory formula [18, 76]
@@
obtained from Eq. (36), which reports $\eta_{\phi}=\eta_{\Gamma}$. Indeed, by replacing $N=25$ in the above large-$N$ formula predicts $y_{\Gamma}\approx 0.459$. This result provides additional evidence of the charged nature of the FP controlling the critical behavior along the CH transition line for sufficiently large $N$.
Figure 6: Data for $N=25$ obtained by varying $J$ across the CH transition line, keeping $\kappa=0.4$ fixed. (Left)(Adapted from Ref. [87]) Plot of $\xi_{Q}/L$, where $\xi_{Q}$ is the second-moment correlation length determined from the two-point correlation function of $Q_{\bm{x}}^{ab}$, defined in Eq. (57); in the main panel data are plotted versus of $(J-J_{c})L^{1/\nu}$, with $J_{c}=0.2955$ and $\nu=0.802$ (the best-fit estimate), in the inset versus $J$. (Right) (Adapted from Ref. [91]) Plot of the Binder parameter $U_{z}$ of the scalar field ${\bm{z}}_{\bm{x}}$ computed in the Lorenz gauge as a function of the ratio $\xi_{Q}/L$, where $\xi_{Q}$ is defined above (see A.3.3 for a discussion of the expected FSS behavior). The excellent data collapse signals that both the gauge-invariant correlation function $G_{Q}$ and the correlation function $G_{z}$ (computed in the Lorenz gauge) are critical along the CH transition line.
@@
Finally, let us discuss the gauge-field correlations. As discussed in Sec. 4.1.2, AHFT predicts $\eta_{A}=1$ at the charged FP. Numerical results along the CH line are in good agreement with this prediction. It is interesting to observe that the result $\eta_{A}=1$ also holds at IXY transitions, see Sec. 5.2.2. However, the IXY value cannot be explained by using the AHFT result, since the AHFT is apparently only relevant for the charged CH transitions and not for the topological IXY transition.
@@ 5.3.4 The Molecular-Higgs transition line
The MH line separates two ordered phases in which the SU($N$) symmetry is broken to U($N-1$[64]. MH transitions are characterized by the condensation of the charged nonlocal excitations and of the gauge modes. As shown in Ref.  [93] the critical behavior along the MH line is the same as along the CH line in the one-component LAH model, see Fig. 2. Therefore, along the MH line one expects IXY transitions. This is confirmed by the analyses of the energy cumulants and of the charged correlation functions (computed in the Lorenz gauge) reported in Ref. [93]. It is important to stress that the decoupling of the scalar modes is not obvious, because of the presence of $22N-$ 2 massless Goldstone bosons related with the spontaneous breaking of the SU($N$) symmetry [64]. These massless excitations turn out not to play any role in the critical behavior along the MH transitions.
As an example, in Fig. 3 we report a scaling plot for the third energy cumulant for $N=2$ and $N=25$. Their FSS behavior is the same as at transitions in the $N=1$ model along the CH line, demonstrating that the transitions belong to the IXY universality class for any $N$. Charged excitations deconfine at the MH transitions. The RG dimension $y_{\Gamma}$ of the charged operator ${\bm{\Gamma}}_{\bm{x}}$, estimated from the critical behavior of the vector variable ${\bm{z}}_{\bm{x}}$ in the Lorenz gauge (see Sec. 5.1.2), turns out [93] to be independent of $N$ and perfectly consistent with the estimate (61) obtained along the CH line of the one-component LAH model, see Fig. 4. Finally, also the critical behavior of the gauge correlations (again computed in the Lorenz gauge) is consistent with an IXY critical behavior. For instance, one obtains $\eta_{A}=1$ also along the MH line [93].
@@ 5.4 $N$-component SO($N$)-invariant lattice Abelian Higgs model
$H_{O}({\bm{A}},{\bm{z}})=\frac{\kappa}{2}\sum_{{\bm{x}},\mu>\nu}F_{{\bm{x}},\mu\nu}^{2}-2NJ\sum_{{\bm{x}},\mu}{\rm Re}\,(\lambda_{{\bm{x}},\mu}\bar{\bm{z}}_{\bm{x}}\cdot{\bm{z}}_{{\bm{x}}+\hat{\mu}})+v\,\sum_{\bm{x}}|{\bm{z}}_{\bm{x}}\cdot{\bm{z}}_{\bm{x}}|^{2}.$ (63)
@@
One can easily check that, after relaxing the unit-length constraint for the scalar field, this Hamiltonian corresponds to the formal continuum limit of the SO($N$)-symmetric AHFT with $\kappa=g^{-2}$,
@@
The phase diagrams at fixed values of $v$ are analogous to that of the SU($N$)-symmetric LAH model, see the right panel of Fig. 2. The qualitative features of the three phases are the same, the only differences being the symmetry of the two ordered phases and the nature of the transitions that depend on the specific symmetry-breaking pattern. These two properties are determined by the specific form of the scalar potential and, in particular, by the sign of the Hamiltonian parameter $v$. The critical behaviors along the various transition lines have been investigated in Refs. [282, 296]. Here we summarize the main results.
@@
The MH transitions are expected to always belong to the IXY universality class, as in SU($N$)-invariant LAH models. This is due to the fact that MH transitions are topological transitions driven by the gauge modes, and scalar fields play no role at the transition, as discussed in Sec. 5.3.4. On the other hand, since the symmetry of the ordered phases depends on the parameter $v$ (the ordered phases are symmetric under SO($N-2$) and ${\rm SO}(N-1)$ transformations for $v>0$ and $v<0$, respectively), the nature of the CH and CM transitions depends on the sign of $v$. Correspondingly, also the relevant order parameter at the transition depends on the sign of the parameter. In particular, the transitions can be specified by the condensation of the lattice bilinear operators $R_{L,{\bm{x}}}^{ab}$ and $T_{L,{\bm{x}}}^{ab}$ defined as their field-theoretical counterparts, see Eq. (40).
Along the CH line one expects charged continuous transitions for a large number $N$ of components. The continuous or first-order nature of the transitions is determined by the presence or absence of the stable FP in the corresponding AHFT. As discussed in Sec. 4.2.1, a FP exists only for $N>N_{o}^{*}(d=3)$. Moreover, the attraction domain of the stable FP is limited to the parameter region $v>0$, i.e., to systems in which the ordered phase is SO($N-2$)-symmetric. Therefore, for $N<N_{o}^{*}(d=3)$ we expect only first-order transitions along the CH line. For $N>N_{o}^{*}(d=3)$, if the Higgs phase is symmetric under SO$(N-1)$ transformations, then the CH line is again a line of first-order transitions. Instead, if the Higgs phase is SO$(N-2)$-symmetric, i.e., $v$ is positive, then the CH transitions can be continuous, in the universality class associated with the stable charged FP of the AHFT (37).
@@
Along the CM line gauge modes do not develop critical correlations, thus CM transitions admit a conventional LGW effective description in terms of a gauge-invariant order parameter [282]. For negative $v$, transitions are driven by the condensation of $R_{\bm{x}}^{ab}$, defined in Eq. (40). A straightforward LGW analysis predicts first-order transitions for any $N\geq 3$. For $N=2$ continuous transitions in the XY universality class are possible. For positive $v$, the critical behavior can be effectively described by the O($N$)-symmetric LGW theory for an antisymmetric rank-two tensor, which is the coarse-grained analogue of $T_{\bm{x}}^{ab}$, see Eq. (40). This allows one to predict that, for $N=2$, continuous CM transitions belong to the Ising universality class for all positive values of $v$. For $N=3$, instead, one expects the existence of a tricritical value $v^{*}>0$, such that the transition is of first order for $v<v^{*}$ and is continuous, in the Heisenberg universality class, for $v>v^{*}$. The existence of a tricritical value is due to the first-order nature of the CM transitions in SU(3) invariant AH models, i.e., for $v=0$. The same behavior, with critical Heisenberg transitions for large values of $v$, occurs also for $N=4$. For $N\geq 5$ transitions are always of first order. A numerical check of these predictions is reported in Ref. [296].
@@ 6 Lattice Abelian Higgs models with compact U(1) gauge variables
In this section we consider an alternative lattice discretization of the AHFT (18) based on compact gauge variables. In compact formulations there are topological features that are absent in noncompact formulations, like the presence of monopole configurations and a nontrivial dependence on the charge $Q$ of the scalar field. As we shall see, these features give rise to notable differences with respect to the noncompact lattice formulation considered in Sec. 5. In particular, the relation with the continuum AHFT (18) turns out to be much less straightforward.
@@ 6.1 The compact formulation of lattice Abelian Higgs models
In the compact formulation on a cubic lattice the gauge fields are complex variables $\lambda_{{\bm{x}},\mu}$ satisfying $|\lambda_{{\bm{x}},\mu}|=1$, i.e., elements of the U(1) group, associated with the lattice links. The compact LAH model with $N$-component scalar fields ${\bm{z}}_{\bm{x}}$ of integer charge $Q$ is defined by the partition function and lattice Hamiltonian
@@
where the two sums in $H_{c}$ run over all plaquettes and links of the cubic lattice, respectively. The parameter $\kappa\geq 0$ plays the role of inverse (square) gauge coupling, as can be seen by taking the naive continuum limit in which the link variables are close to one. The compact formulation is well defined for all boundary conditions, since the Hamiltonian is bounded and the configuration space is compact. While in the noncompact case, the gauge group is the noncompact additive group of real numbers, in the compact case the model is invariant under U(1) gauge transformations. The nontrivial topology of the U(1) group allows one to define an integer charge $Q$ and consider scalar fields that transform under a charge-$Q$ representation of the U(1) group. The charge-$Q$ model is invariant under the U(1) gauge transformations
@@
Unlike what happens in noncompact formulations of LAH models, the integer charge $Q$ cannot be absorbed in the definition of the gauge field [297] and is, therefore, an additional independent Hamiltonian parameter. As we shall see, the phase diagram of compact LAH models crucially depends on the charge $Q$. In particular, the $\kappa$-$J$ phase diagram of models with $Q=1$ differs from that of models with $Q\geq 2$. Moreover, as in the noncompact case, the behavior for $N=1$ is qualitatively different from that obtained for $N\geq 2$.
@@
The unitary gauge is not complete and leaves a residual invariance under local ${\mathbb{Z}}_{Q}$ gauge transformations, i.e., under the transformations $\lambda_{{\bm{x}},\mu}=V_{\bm{x}}\lambda_{{\bm{x}},\mu}\bar{V}_{{\bm{x}}+\hat{\mu}}$, where $V_{\bm{x}}$ are complex numbers satisfying $V_{\bm{x}}^{Q}=1$, i.e., elements of the ${\mathbb{Z}}_{Q}$ group. Topological transitions can thus be present, cf. Sec. 3.5.3, since for $Q\geq 2$ the model with Hamiltonian (66) is a soft version of the ${\mathbb{Z}}_{Q}$ gauge theory defined in Sec. 3.5.2, which is obtained in the limit $J\to\infty$. Indeed, for $J\to\infty$, we should have $\lambda_{{\bm{x}},\mu}^{Q}=1$ on all links, which implies that the fields $\lambda_{{\bm{x}},\mu}$ are elements of the discrete ${\mathbb{Z}}_{Q}$ group. For $Q=1$ the limit is obviously trivial.
A line of topological transition is present, when $Q\geq 2$ also in the broken phase of the $N>1$ model: the minima of the potential correspond to (in the thermodynamic limit and up to a gauge transformation) to ${\bm{z}}_{\bm{x}}={\bm{z}}_{{\bm{x}}+\hat{\mu}}$ and $\lambda_{{\bm{x}},\mu}^{Q}=1$, hence a residual $\mathbb{Z}_{Q}$ gauge symmetry is present also in this case.
6.2 The compact lattice Abelian Higgs model with one-component scalar matter
Figure 8: Sketch of the $J$-$\kappa$ phase diagram of the 3D one-component CLAH model, in which a compact U(1) gauge field is coupled with a (unit-length) complex scalar field of charge $Q\geq 2$. A confined phase (for small $J$ or small $\kappa$) and a deconfined phase (for large $J$ and $\kappa$) are present, separated by a deconfinement transition line. For $\kappa\to\infty$ and $J\to\infty$ the model reduces to the XY vector model and to the lattice ${\mathbb{Z}}_{Q}$ gauge model with Wilson action, respectively.

### v3

@@ 5.3.3 The Coulomb-Higgs transition line
At the continuous CH transitions scalar, charged, and gauge modes are all critical. As an example, in Fig. 6 we report the results [87, 91] of FSS analyses of numerical data (see A.3 for a summary of the relevant FSS relations used in the numerical analysis), that provide evidence of a continuous transition for $N=25$ with critical scalar modes. For instance, the left panel of Fig. 6 shows the scaling behavior of the correlation length $\xi_{Q}$ computed in terms of the correlation function defined in Eq. (61), while the right panel shows the scaling behavior of the Binder parameter of the scalar field ${\bm{z}}_{\bm{x}}$ versus the RG invariant ratio $\xi_{Q}/L$, computed in the Lorenz gauge (51) (in this gauge the correlations of the local field ${\bm{z}}_{\bm{x}}$ are equal to those of the nonlocal gauge-invariant charged operator ${\bm{\Gamma}}_{\bm{x}}$ defined in Eq. (54), see Sec. 5.1.2). The excellent quality of the data collapse confirms that the transition is continuous and that correlations of both $Q_{\bm{x}}$ and ${\bm{\Gamma}}_{\bm{x}}$ are critical. Similar results have been obtained for the gauge-field correlation functions.
It is important to note that the choice of the hard Lorenz gauge fixing (51) turns out to be crucial to observe the condensation of the scalar field ${\bm{z}}_{\bm{x}}$. If, instead of imposing the Lorenz constraint (51), we add the gauge-fixing term (52) to the Hamiltonian (soft Lorenz gauge fixing), then scalar-field correlations are always short ranged and the gauge-field correlations show the typical Coulomb behavior, for any positive value of the gauge parameter $\zeta$, irrespective of the phase one is considering. As already noted in Sec. 4.1.3, the hard Lorenz gauge fixing, which is obtained in the limit $\zeta\to 0$, is also singled out by the AHFT. Since $\zeta=0$ is an unstable FP of the RG flow, see Eq. (34), if we start the RG flow from any finite $\zeta>0$, then $\zeta$ increases towards infinity, so the nongauge-invariant modes become unbounded. The instability of the $\zeta=0$ FP, implies a nontrivial crossover behavior for small values of $\zeta$, characterized by a length scale $\xi_{\zeta}\sim\zeta^{-1}$. This prediction has been verified numerically in Ref. [91] for the noncompact LAH model.
@@
In Fig. 7 the numerical estimates of the exponent $\nu$ and $\eta_{q}$ are compared with the $1/N$ expressions reported in Sec. 4.1.4. The agreement is very good, thus providing a stringent check that the AHFT provides the correct effective description of the continuous transitions along the CH line. The same check has been performed for the exponent $\eta_{\Gamma}$, which has been determined by analyzing the susceptibility $\chi_{z}$ of the scalar field ${\bm{z}}_{\bm{x}}$ computed in the Lorenz gauge (51). Ref. [91] estimated $\eta_{\Gamma}$ for $N=25$, which, in turn provides an estimate of the RG dimension $y_{\Gamma}=(1+\eta_{\Gamma})/2$ of ${\bm{\Gamma}}_{\bm{x}}$. The result $y_{\Gamma}=0.4655(5)$ compares satisfactorily with the large-$N$ field-theory formula [18, 76]
@@
obtained from Eq. (38), which reports $\eta_{\phi}$ in the hard Lorenz gauge (in this case $\eta_{\phi}=\eta_{\Gamma}$). Indeed, for $N=25$ the large-$N$ formula (65) predicts $y_{\Gamma}\approx 0.459$. This result provides additional evidence of the charged nature of the FP controlling the critical behavior along the CH transition line for sufficiently large $N$.
Figure 6: Data for $N=25$ obtained by varying $J$ across the CH transition line, keeping $\kappa=0.4$ fixed. (Left)(Adapted from Ref. [87]) Plot of $\xi_{Q}/L$, where $\xi_{Q}$ is the second-moment correlation length determined from the two-point correlation function of $Q_{\bm{x}}^{ab}$, defined in Eq. (60); in the main panel data are plotted versus of $(J-J_{c})L^{1/\nu}$, with $J_{c}=0.2955$ and $\nu=0.802$ (the best-fit estimate), in the inset versus $J$. (Right) (Adapted from Ref. [91]) Plot of the Binder parameter $U_{z}$ of the scalar field ${\bm{z}}_{\bm{x}}$ computed in the Lorenz gauge as a function of the ratio $\xi_{Q}/L$, where $\xi_{Q}$ is defined above (see A.3.4 for a discussion of the expected FSS behavior). The excellent data collapse signals that both the gauge-invariant correlation function $G_{Q}$ and the correlation function $G_{z}$ (computed in the Lorenz gauge) are critical along the CH transition line.
@@
Finally, let us discuss the gauge-field correlations. As discussed in Sec. 4.1.2, AHFT predicts $\eta_{A}=1$ at the charged FP. Numerical results along the CH line are in good agreement with this prediction. It is interesting to observe that the result $\eta_{A}=1$ also holds at IXY transitions, see Sec. 5.2.2. However, the IXY value cannot be explained by using the AHFT result, since the AHFT is apparently only relevant for the charged CH transitions and not for the topological IXY transition (see, however, the discussion reported at the end of Sec. 5.2.2).
@@ 5.3.4 The Molecular-Higgs transition line
The MH line separates two ordered phases in which the SU($N$) symmetry is broken to U($N-1$[64]. MH transitions are characterized by the condensation of the charged nonlocal excitations and of the gauge modes. As shown in Ref. [93] the critical behavior along the MH line is the same as that along the CH line in the one-component LAH model, see Fig. 2. Therefore, along the MH line one expects IXY transitions. This is confirmed by the analyses of the energy cumulants and of the charged correlation functions (computed in the Lorenz gauge) reported in Ref. [93]. It is important to stress that the decoupling of the scalar modes is not obvious, because of the presence of $22N-$ 2 massless Goldstone bosons related with the spontaneous breaking of the SU($N$) symmetry [64]. These massless excitations turn out not to play any role in the critical behavior along the MH transitions.
As an example, in Fig. 3 we report a scaling plot of the third energy cumulant for $N=2$ and $N=25$. Their FSS behavior is the same as that at the CH transitions in the $N=1$ model, demonstrating that the transitions belong to the IXY universality class for any $N$. Charged excitations deconfine at the MH transitions. The RG dimension $y_{\Gamma}$ of the charged operator ${\bm{\Gamma}}_{\bm{x}}$, estimated from the critical behavior of the vector variable ${\bm{z}}_{\bm{x}}$ in the Lorenz gauge (see Sec. 5.1.2), turns out [93] to be independent of $N$ and perfectly consistent with the estimate (64) obtained along the CH line of the one-component LAH model, see Fig. 4. Finally, also the critical behavior of the gauge correlations (again computed in the Lorenz gauge) is consistent with an IXY critical behavior. For instance, one obtains $\eta_{A}=1$ also along the MH line [93].
@@ 5.4 $N$-component SO($N$)-invariant lattice Abelian Higgs model
$H_{O}({\bm{A}},{\bm{z}})=\frac{\kappa}{2}\sum_{{\bm{x}},\mu>\nu}F_{{\bm{x}},\mu\nu}^{2}-2NJ\sum_{{\bm{x}},\mu}{\rm Re}\,(\lambda_{{\bm{x}},\mu}\bar{\bm{z}}_{\bm{x}}\cdot{\bm{z}}_{{\bm{x}}+\hat{\mu}})+v\,\sum_{\bm{x}}|{\bm{z}}_{\bm{x}}\cdot{\bm{z}}_{\bm{x}}|^{2},$ (66)
@@
where again we consider the unit-length constraint $\bar{\bm{z}}_{\bm{x}}\cdot{\bm{z}}_{\bm{x}}=1$. One can easily check that, after relaxing the unit-length constraint for the scalar field, this Hamiltonian corresponds to the formal continuum limit of the SO($N$)-symmetric AHFT with $\kappa=g^{-2}$.
@@
The phase diagram for fixed values of $v$ is analogous to that of the SU($N$)-symmetric LAH model, see the right panel of Fig. 2. The qualitative features of the three phases are the same, the only differences being the symmetry of the two ordered phases and the nature of the transitions that depend on the specific symmetry-breaking pattern. These two properties are determined by the specific form of the scalar potential and, in particular, by the sign of the Hamiltonian parameter $v$. The critical behaviors along the various transition lines have been investigated in Refs. [314, 324]. Here we summarize the main results.
@@
The MH transitions are expected to always belong to the IXY universality class, as in SU($N$)-invariant LAH models. This is due to the fact that MH transitions are topological transitions driven by the gauge modes, and scalar fields play no role at the transition, as discussed in Sec. 5.3.4. On the other hand, since the symmetry of the ordered phases depends on the parameter $v$ (the ordered phases are symmetric under SO($N-2$) and ${\rm SO}(N-1)$ transformations for $v>0$ and $v<0$, respectively), the nature of the CH and CM transitions depends on the sign of $v$. Correspondingly, also the relevant order parameter at the transition depends on the sign of the parameter. In particular, the transitions can be characterized by the condensation of the lattice bilinear operators $R_{L,{\bm{x}}}^{ab}$ and $T_{L,{\bm{x}}}^{ab}$ defined as their field-theoretical counterparts, see Eq. (42).
[…]
@@
Along the CM line gauge modes do not develop critical correlations, thus CM transitions admit a conventional LGW effective description in terms of a gauge-invariant order parameter [314]. For negative $v$, transitions are driven by the condensation of $R_{\bm{x}}^{ab}$, defined in Eq. (42). A straightforward LGW analysis predicts first-order transitions for any $N\geq 3$. For $N=2$ continuous transitions in the XY universality class are possible. For positive $v$, the critical behavior can be effectively described by the O($N$)-symmetric LGW theory for an antisymmetric rank-two tensor, which is the coarse-grained analogue of $T_{\bm{x}}^{ab}$ defined in Eq. (42). This allows us to predict that, for $N=2$, continuous CM transitions belong to the Ising universality class for all positive values of $v$. For $N=3$, instead, one expects the existence of a tricritical value $v^{*}>0$, such that the transition is of first order for $v<v^{*}$ and is continuous, in the Heisenberg universality class, for $v>v^{*}$. The existence of a tricritical value is due to the first-order nature of the CM transitions in SU(3) invariant AH models, i.e., for $v=0$. The same behavior, with critical Heisenberg transitions for large values of $v$, occurs also for $N=4$. For $N\geq 5$ transitions are always of first order. A numerical check of these predictions is reported in Ref. [324].
@@ 6 Lattice Abelian Higgs models with compact U(1) gauge variables
In this section we consider an alternative lattice discretization of the AHFT (18) based on compact gauge variables. In compact formulations there are topological features that are absent in noncompact formulations, like the presence of monopole configurations and a nontrivial dependence on the charge $Q$ of the scalar field. As we shall see, these features give rise to notable differences with respect to the noncompact lattice formulation considered in Sec. 5. In particular, the relation with the continuum AHFT (18) turns out to be much less straightforward. The nontrivial relation between the LAH models with noncompact gauge fields and the compact LAH models with higher-charge scalar fields are discussed at the end of the section.
Like noncompact LAH models, we must distinguish between one-component and multi-component scalar models, which show qualitatively different phase diagrams. As before, this is related to the presence of the global SU($N$) symmetry for $N\geq 2$. Moreover, the behavior of unit-charge ($Q=1$) models differs substantially from that of models in which the scalar fields have charge $Q\geq 2$. Indeed, for $Q\geq 2$ the models also undergo topological transitions analogous to those in lattice ${\mathbb{Z}}_{Q}$ models (thus depending on the charge $Q$), see Sec. 3.5.2.
@@ 6.1 The compact formulation of lattice Abelian Higgs models
In the compact formulation the gauge fields are complex variables $\lambda_{{\bm{x}},\mu}$ satisfying $|\lambda_{{\bm{x}},\mu}|=1$, i.e., elements of the U(1) group, associated with the lattice links. On a cubic lattice, the compact LAH model with $N$-component scalar fields ${\bm{z}}_{\bm{x}}$ of integer charge $Q$ is defined by the partition function and lattice Hamiltonian
@@
where the two sums in $H_{c}$ run over all plaquettes and links of the cubic lattice, respectively, and the site variables satisfy the unit-length condition $\bar{\bm{z}}_{\bm{x}}\cdot{\bm{z}}_{{\bm{x}}}=1$. The parameter $\kappa\geq 0$ plays the role of inverse (square) gauge coupling, as can be seen by taking the naive continuum limit in which the link variables are close to one. The compact formulation is well defined for all boundary conditions, since the Hamiltonian is bounded and the configuration space is compact. While in the noncompact case, the gauge group is the noncompact additive group of real numbers, in the compact case the model is invariant under U(1) gauge transformations. The nontrivial topology of the U(1) group allows one to define an integer charge $Q$ and consider scalar fields that transform under a charge-$Q$ representation of the U(1) group. The charge-$Q$ model is invariant under the U(1) gauge transformations
@@
Unlike what happens in noncompact formulations of LAH models, the integer charge $Q$ cannot be absorbed in the definition of the gauge field [325]. Therefore, it should be considered as an additional independent Hamiltonian parameter. As we shall see, the phase diagram of compact LAH models crucially depends on the charge $Q$. In particular, the $\kappa$-$J$ phase diagram of models with $Q=1$ differs from that of models with $Q\geq 2$. Moreover, as in the noncompact case, the behavior for $N=1$ is qualitatively different from that obtained for $N\geq 2$.
@@
The unitary gauge is not complete and leaves a residual invariance under local ${\mathbb{Z}}_{Q}$ gauge transformations, i.e., under the transformations $\lambda_{{\bm{x}},\mu}=V_{\bm{x}}\lambda_{{\bm{x}},\mu}\bar{V}_{{\bm{x}}+\hat{\mu}}$, where $V_{\bm{x}}$ are complex numbers satisfying $V_{\bm{x}}^{Q}=1$, i.e., elements of the ${\mathbb{Z}}_{Q}$ group. For $Q\geq 2$ topological transitions can be present, as discussed in Sec. 3.5.3, since the model with Hamiltonian (69) is a soft version of the ${\mathbb{Z}}_{Q}$ gauge theory defined in Sec. 3.5.2, which is obtained in the limit $J\to\infty$. Indeed, for $J\to\infty$, we should have $\lambda_{{\bm{x}},\mu}^{Q}=1$ on all links, which implies that the fields $\lambda_{{\bm{x}},\mu}$ are elements of the discrete ${\mathbb{Z}}_{Q}$ group. For $Q=1$ the limit is obviously trivial.
6.2 The compact lattice Abelian Higgs model with one-component scalar fields
Figure 8: Sketch of the $\kappa$-$J$ phase diagram of the 3D one-component CLAH model, in which a compact U(1) gauge field is coupled with a (unit-length) complex scalar field of charge $Q\geq 2$. A confined phase (for small $J$ or small $\kappa$) and a deconfined phase (for large $J$ and $\kappa$) are present, separated by a deconfinement transition line. For $\kappa\to\infty$ and $J\to\infty$ the model reduces to the XY vector model and to the lattice ${\mathbb{Z}}_{Q}$ gauge model with Wilson action, respectively.